OpenStudy (anonymous):

What is the value of x in the equation 4(2x + 10) = 0?

OpenStudy (anonymous):

im back lol

OpenStudy (anonymous):

lol dude your my number 1 helper.

OpenStudy (anonymous):

x=-5

alones (alones):

4(2x+10)=0 simplify both sides of the equation. 4(2x+10)=/0 8x+40=0 subtract 40 from both sides. 8x+40−40=0−40 8x=−40 the divide both sides by 8. 8x/8=−408 x=−5

alones (alones):

@clanger555 but you didn't explain how you got the answer

OpenStudy (anonymous):

i was gonna to

OpenStudy (anonymous):

4(2x+10)=0 2x+10=0 2x=-10 x= -5

OpenStudy (anonymous):

8x+40-40=0-40 8x=-40 \frac{8x}{8}=\frac{-40}{8} x=-5

alones (alones):

clanger you first explain like @ajith then give the answer -.-

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thankks guys